One solution is 14% acid and another is 19% acid. How much of each solution is needed to make 100 mL solution that is 15% acid?
Answer by robertb(5830)

0.14x + 0.19(100 - x) = 0.15*100
==> 0.14x + 19 - 0.19x = 15
==>-0.05x = -4
==> x = 80 ml of the 14% acid solution, and
100 - x = 20 ml of the 19% acid solution.